The decomposition of ozone inane stranosphere can occ by the two sol which species is an intermediate in this mechanism? species at equilibrium are as follows: given that the o 60 moln 0111 027 movi.. 1- 0.85 mol/1. [12 B. 0.22 D. 0.19 E. 1.6 x 102 22. Calculate Kp for the reaction 2Noctogo 2NOG) Cla(e) at 40o C if Ke at 400 Cfor this reaction is 2.1 x 10 2 A. 2.1 x 10 B. 1.7 x 10-3 C. 0.70 D. 1.2 3.8 x 10-4
Answer #1

20) the species which is formed in one step and is consumed in some other step is called as intermediate

so in the given reaction BrO is the intermediate

21) Kc for the reaction will be

Kc = [H2] [I2] / [HI]2

Kc = 0.27 X 0.6 / (0.85)^1/2 = 0.224 = 0.22

22) We know that

Kp = Kc (RT)Δn

Where

Δn = number of moles of gaseous products - number of moles of gaseous reactants

Here Δn = 3-2 =1

R= gas constant= 0.0821 L atm /mol K

T= temperature = 400 C = 400 +273 K = 673K

Kp = 2.1 X 10^-2 (0.0821 X 673) = 1.160 = 1.2